Calculate the pH for each of the following cases in the titration of 50.0 mL of 0.230 M HClO(aq) with 0.230 M KOH(aq). The ionization constant for HClO can be found here.

(a) before addition of any KOH

(b) after addition of 25.0 mL of KOH

(c) after addition of 35.0 mL of KOH

(d) after addition of 50.0 mL of KOH

(e) after addition of 60.0 mL of KOH
Answer #1

), DATE PAGE Helo - coce it f co a ce ke a cox ca a C- ca [ut] a G = cх ke Shox 4+1 = √2.9810-8 x 0.93 +] = 0.667 X10 [+] = 0DATE PAGE This is a buffore Belution H 2 pro & log salt? leida Ha -log (2.9 410-8) + log soff pH = -log 2.9 + 8 2 7.53 To & -DATE PAGE Holo t somL kok s cooke + 0.23M som lo ke & Ho H₂O 0.23M 11.5mmule 11. simmel 11.5mmal This is a hydrolysis whept.